# SVD-LLM: Truncation-aware Singular Value Decomposition for Large Language Model Compression

### v2

@@ 4.5 Benefits to other LLM Compression Methods
SVD-LLM is orthogonal to other LLM compression methods including quantization and parameter pruning. In this experiment, we add the techniques proposed in SVD-LLM into representative quantization and parameter pruning-based LLM compression methods that are widely recognized by the community to examine how SVD-LLM could further enhance their performance.
Integrate SVD-LLM with Quantization. We select GPTQ (Frantar et al., 2022) as the quantization method. Specifically, we compress LLaMA-7B by GPTQ-4bit combined with SVD-LLM, and compare the compressed model against LLaMA-7B compressed by GPTQ-3bit. As shown in Table 8, combining GPTQ-4bit with SVD-LLM achieves a perplexity that is $18\%$ lower than GPTQ-3bit even with a smaller memory footprint (2.1 GB vs. 2.8 GB). This result demonstrates that compared to directly quantizing using smaller number of bits, GPTQ is able to achieve better compression performance with the help of SVD-LLM.
@@
Integrate SVD-LLM with Parameter Pruning. We select LLM-Pruner (Ma et al., 2023) as the parameter pruning method. Specifically, we compress LLaMA-7B by LLM-Pruner under 30% compression ratio combined with SVD-LLM, and compare the compressed model against LLaMA-7B compressed by LLM-Pruner under 40% compression ratio. As shown in Table 9, LLM-Pruner achieves better compression performance when used in conjunction with SVD-LLM. In particular, with the same memory footprint of 8.8 GB, combining LLM-Pruner under 30% compression ratio with SVD-LLM achieves a perplexity that is $13\%$ lower than LLM-Pruner under 40% compression ratio.
Table 9: Perplexity ($\downarrow$) of LLaMA-7B compressed by LLM-Pruner w/ and w/o SVD-LLM on WikiText-2. The relative performance gain of combined compression compared to LLM-Pruner under 40% compression ratio is marked in green color inside bracket.
@@
4.6 KV Cache Compression with SVD-LLM
SVD-LLM is able to not only compress LLMs but also compress the runtime KV cache at the same time. Specifically, instead of keeping the original intermediate state matrix $m=WX$ with shape $M\times L$ inside the KV cache, after decomposing and compressing $W$ into $W_{u}W_{v}^{T}$, SVD-LLM only needs to store $m^{\prime}=W_{v}^{T}X$ with shape $r\times L$. Therefore, the size of KV cache can be compressed to $\frac{r}{M}$ of the original one. Moreover, since $W_{u}$ is already stored as the weight matrix in the decomposed LLM, the original intermediate state matrix can still be recovered by $m=W_{u}m^{\prime}$ without accuracy drop. Therefore, SVD-LLM provides a unified solution that combines model compression and KV cache compression into a single process. This is different from existing quantization or parameter pruning-based LLM compression methods that need to be combined with other techniques for compressing both weights and KV cache.
In our last experiment, we evaluate this additional benefit on KV cache compression brought by SVD-LLM. This is a new avenue given that KV cache compression has not been evaluated in previous LLM compression studies. Specifically, we measure the peak memory footprint during inference at runtime when generating 128 tokens with batch size of 32 using LLaMA-7B compressed by SVD-LLM under different compression ratios w/ and w/o considering KV cache compression. The results are illustrated in Figure 5 where the difference between blue and yellow bars marked in red represents the reduced footprint of KV cache. As shown, SVD-LLM is able to effectively reduce the footprint of KV cache, and therefore, the peak memory during inference at runtime across all the compression ratios.
Figure 5: Peak memory to generate 128 tokens with batch size of 32 using LLaMA-7B compressed by SVD-LLM under different compression ratios w/ and w/o KV-cache compression. The difference between blue and yellow bars marked in red indicates the reduced footprint of KV cache.
@@ 5 Conclusion
In this paper, we presented SVD-LLM, a SVD-based LLM compression method. SVD-LLM proposes a novel truncation-aware data whitening strategy to guide which singular values to be truncated with minimal compression loss. It also introduces a layer-wise closed-form model parameter update scheme to compensate for accuracy degradation under high compression ratios. We have demonstrated the effectiveness of SVD-LLM on 11 datasets and seven models from three LLM families at four scales and have shown its superiority over state-of-the-arts. We also show its effectiveness in further enhancing the performance of other LLM compression methods as well as KV cache reduction.
6 Broader Impacts
We present SVD-LLM, a post-training SVD technique that effectively compresses large language models (LLMs) while maintaining accuracy and robustness. Our method achieves state-of-the-art performance, particularly when dealing with high compression ratios. We demonstrate the value and potential of SVD compression for LLMs by investigating two aspects: (1) the ability of SVD compression to reduce the size of the KV cache, resulting in savings in LLM runtime memory without additionally sacrificing accuracy, and (2) the role of SVD compression in enhancing the compression performance of other LLM compression techniques. Our work contributes to the advancement of compression methods that improve the efficiency and deployability of LLMs for real-world applications.
@@ A.1 The compression error of ASVD
which is still a complex function that involves the activation $X$, the scaling matrix $S_{0}$, the singular vector $v^{\prime}_{i}$ and the singular value $\sigma^{\prime}_{i}$. As a result compression error is not directly related to the singular value and the conventional SVD compression by truncating the smallest singular values may lead to suboptimal compression error.

### v3

@@ 4.3 Ablation Study
Modular Sensitivity Study: We conduct ablation studies to evaluate the separate contributions of the two key components (truncation-aware data whitening and layer-wise closed-form update) of SVD-LLM. Let SVD-LLM (W) denote the version of SVD-LLM with truncation-aware data whitening only; SVD-LLM (U) denote the version of SVD-LLM with layer-wise closed-form update only; and SVD-LLM (W+U) denote the version of SVD-LLM with both truncation-aware data whitening and layer-wise closed-form update. The results are shown in Section 4.3. We have three observations. (1) Both SVD-LLM (W) and SVD-LLM (U) consistently outperform ASVD across all the compression ratios. Notably, when the compression ratio is at and above 40%, both variants reduce the perplexity by more than 99% compared to ASVD. (2) Under 20% and 30% compression ratios, SVD-LLM (W) achieves the lowest perplexity compared to SVD-LLM (U) and SVD-LLM (W+U). (3) Under 40%, 50% and 60% compression ratios, SVD-LLM (W+U) achieves the lowest perplexity compared to SVD-LLM (W) and SVD-LLM (U), highlighting the importance of combining both truncation-aware data whitening and layer-wise closed-form update when compression ratio goes high.
Calibration Data Analysis: We next analyze the impact of calibration data used for both truncation-aware data whitening and layer-wise closed-form update on the compression performance. LABEL:fig:cali and Section 4.2 summarize the performance of compressed LLaMA-7B when changing three key characteristics of the calibration data, the number of the calibration data, the seed used to randomly sample the calibration data, and the data set from which the calibration data is sampled. As shown, changing any of the three characteristics only causes a tiny disturbance of less than 15% to the final performance, demonstrating that SVD-LLM is less sensitive to the design of the calibration data to compress LLM.
@@ 4.4 Benefits to other LLM Compression Methods
SVD-LLM is orthogonal to other LLM compression methods including quantization and parameter pruning. In this experiment, we combine SVD-LLM with quantization and parameter pruning-based LLM compression methods that are widely recognized by the community to examine how SVD-LLM could further enhance their performance.
Integrate SVD-LLM with Quantization. We select GPTQ [6] as the quantization method. Specifically, we compress LLaMA-7B by GPTQ-4bit combined with SVD-LLM, and compare the compressed model against LLaMA-7B compressed by GPTQ-3bit. As shown in Section 4.4, combining GPTQ-4bit with SVD-LLM achieves a perplexity that is $18\%$ lower than GPTQ-3bit even with a smaller memory footprint (2.1 GB vs. 2.8 GB). This result demonstrates that compared to directly quantizing using smaller number of bits, GPTQ achieves better compression performance with the help of SVD-LLM.
Integrate SVD-LLM with Parameter Pruning. We select LLM-Pruner [16] as the parameter pruning method. Specifically, we compress LLaMA-7B by LLM-Pruner under 30% compression ratio combined with SVD-LLM, and compare the compressed model against LLaMA-7B compressed by LLM-Pruner under 40% compression ratio. As shown in Table 4.4, LLM-Pruner achieves better compression performance when used in conjunction with SVD-LLM. In particular, with the same memory footprint of 8.8 GB, combining LLM-Pruner under 30% compression ratio with SVD-LLM achieves a perplexity that is $13\%$ lower than LLM-Pruner under 40% compression ratio.
@@
Table 8: Perplexity ($\downarrow$) of LLaMA-7B compressed by LLM-Pruner w/ and w/o SVD-LLM on WikiText-2. The relative performance gain of combined compression compared to LLM-Pruner under 40% compression ratio is marked in green color.
@@
4.5 Benefits of Inference Speedup
SVD-LLM is capable to achieve inference speedup. To demonstrate this advantage, we measure the number of tokens that the original LLaMA-7B and its compressed version by SVD-LLM can generate on average per second with different batch size and sequence length.  LABEL:fig:gpu_throughput show the results on the GPU and CPU. As shown, SVD-LLM consistently ensures an acceleration in the generation speed across all the compression ratios illustrated in the figure. More importantly, this enhancement becomes more significant as the batch size increases and the sequence length decreases, resulting in a maximum speedup of 1.7x on GPU and 1.5x on CPU under the 40% compression ratio, where the model performance remains acceptable according to Section 4.1. These results highlight the effectiveness of SVD-LLM in improving the efficiency of LLM for real-world usage.
Figure 10: Peak memory to generate 128 tokens with batch size of 32 using LLaMA-7B compressed by SVD-LLM under different compression ratios w/ and w/o KV-cache compression. The difference between the blue and yellow bars marked in red indicates the reduced footprint of the KV cache.
4.6 Benefits of KV Cache Compression
SVD-LLM is able to not only compress LLMs but also compress the runtime KV cache at the same time. Specifically, instead of keeping the original intermediate state matrix $m=WX$ with shape $M\times L$ inside the KV cache, after decomposing and compressing $W$ into $W_{u}W_{v}^{T}$, SVD-LLM only needs to store $m^{\prime}=W_{v}^{T}X$ with shape $r\times L$. Therefore, the size of the KV cache can be compressed to $\frac{r}{M}$ of the original. Moreover, since $W_{u}$ is already stored as the weight matrix in the decomposed LLM, the original intermediate state matrix can still be recovered by $m=W_{u}m^{\prime}$ without accuracy drop. Therefore, SVD-LLM provides a unified solution that combines model compression and KV cache compression into a single process. This is different from existing quantization or parameter pruning-based LLM compression methods that need to be combined with other techniques for compressing both weights and KV cache..
In our last experiment, we evaluate this benefit on KV cache compression brought by SVD-LLM. This is a new avenue since KV cache compression has not been evaluated in previous LLM compression studies. Specifically, we measure the peak memory footprint during inference when generating 128 tokens with batch size of 32 using LLaMA-7B compressed by SVD-LLM under different compression ratios w/ and w/o considering KV cache compression. The results are illustrated in Figure 10 where the difference between the blue and yellow bars marked in red represents the reduced footprint of the KV cache. As shown, SVD-LLM is able to effectively reduce the footprint of KV cache. Therefore, the peak memory during inference at runtime across all the compression ratios.
@@ 5 Conclusion
In this paper, we presented SVD-LLM, a SVD-based LLM compression method. SVD-LLM proposes a novel truncation-aware data whitening strategy to guide which singular values to be truncated with minimal compression loss. It also introduces a layer-wise closed-form model parameter update scheme to compensate for accuracy degradation under high compression ratios. We have demonstrated the effectiveness of SVD-LLM on $10$ datasets and seven models from three LLM families at four scales and have shown its superiority over state-of-the-arts. We also show its effectiveness in further enhancing the performance of other LLM compression methods.
@@ A.1 The compression error of ASVD
which is still a complex function that involves the activation $X$, the diagonal matrix $S_{0}$, the singular vector $v^{\prime}_{i}$ and the singular value $\sigma^{\prime}_{i}$. As a result, compression error is not directly related to the singular value, and the conventional SVD compression by truncating the smallest singular values may lead to suboptimal compression error.
A.2 Pseudocode for SVD-LLM
Algorithm 1 shows the pseudocode of SVD-LLM. Before compression, SVD-LLM randomly collects a small amount of sentences as the calibration data $C$, it then runs the truncation-aware data whitening process as shown in Algorithm 2 to obtain the set of whitening matrix $\text{Set}_{S}$ for the weight to compress. After that, it runs the SVD and truncation with $\text{Set}_{S}$ on each weight matrix in the LLM. Instead of directly finishing the whole compression, it stores the decomposed matrices and further utilizes these matrices to run the layer-wise closed-form update as shown in Algorithm 3.
Algorithm 1 Pseudocode for SVD-LLM
1:Input: $M$: Original LLM
2:Output: $M^{\prime}$: Compressed LLM by SVD-LLM
3:procedure SVD-LLM($M$)
4:     Randomly collect several sentences as the calibration data $C$
5:     $\text{Set}_{S}\leftarrow\textsc{Truncation-Aware Data Whitening}(M,C)$
6:     $\text{Set}_{SVD}\leftarrow\emptyset$triangleright Initialize the set of decomposed matrices for the weight to compress
7:     $\text{Set}_{W}\leftarrow M$triangleright Obtain the set of weights in $M$ to compress
8:     for $W$ in $\text{Set}_{W}$ do
9:         $S\leftarrow\text{Set}_{S}(W)$triangleright Extract the whitening matrix of current weight $W$
10:         $U,\Sigma,V\leftarrow\text{SVD}(WS)$triangleright Apply singular value decomposition on $W$
11:         $\Sigma_{1}\leftarrow\text{Trunc.}(\Sigma)$triangleright Truncate the smallest singular values in $\Sigma$
12:         $\text{Set}_{SVD}\leftarrow(U,\Sigma_{1},V)\cap\text{Set}_{SVD}$triangleright Store the decomposed matrices of the weight in the set
13:     end for
14:     $M^{\prime}\leftarrow\textsc{Layer-Wise Closed-Form Update}(M,C,\text{Set}_{S},\text{Set}_{SVD})$
15:     return $M^{\prime}$
16:end procedure
Algorithm 2 Pseudocode for Truncation-Aware Data Whitening
1:Input: $M$: Original LLM
2:Input: $C$: Calibration Data
3:Output: $\text{Set}_{S}$: Set of whitening matrices for the weight to compress in $M$
4:procedure Truncation-Aware Data Whitening($M,C$)
5:     $\text{Set}_{S}\leftarrow\emptyset$triangleright Initialize the set of whitening matrices
6:     $\text{Set}_{W}\leftarrow M$triangleright Obtain the set of weights in $M$ to compress
7:     for $W$ in $\text{Set}_{W}$ do
8:         $X\leftarrow M(W,D)$triangleright Obtain the input activation of the weight matrix $W$
9:         $S\leftarrow\text{Cholesky\_Decomposition}(XX^{T})$triangleright Apply cholesky decomposition on $XX^{T}$
10:         $\text{Set}_{S}\leftarrow S\cup\text{Set}_{S}$triangleright Store the whitening weight matrix in the set
11:     end for
12:     return $\text{Set}_{S}$
13:end procedure
Algorithm 3 Pseudocode for Layer-Wise Closed-Form Update
1:Input: $M$: Original LLM
2:Input: $C$: Calibration Data
3:Input: $\text{Set}_{S}$: Set of whitening matrices for the weight to compress in $M$
4:Input: $\text{Set}_{SVD}$: Set of decomposed matrices for the weight to compress in $M$
5:Output: $M^{\prime}$: Compressed LLM by SVD-LLM
6:procedure Layer-Wise Closed-Form Update($M,C,\text{Set}_{S},\text{Set}_{SVD}$)
7:     $M^{\prime}\leftarrow M$triangleright Initialize $M^{\prime}$ with $M$
8:     $\text{Set}_{L}\leftarrow M^{\prime}$triangleright Obtain the set of encoder and decoder layers in $M^{\prime}$
9:     $X^{\prime}\leftarrow M^{\prime}(C)$triangleright Obtain the input activation of the first layer in $M^{\prime}$
10:     for $L$ in $\text{Set}_{L}$ do
11:         $\text{Set}_{W}\leftarrow L$triangleright Obtain the set of weights in $L$ to compress
12:         for $W$ in $\text{Set}_{W}$ do
13:              $S\leftarrow\text{Set}_{S}(W)$triangleright Extract the whitening matrix of current weight $W$
14:              $U,\Sigma_{1},V\leftarrow\text{Set}_{SVD}(W)$triangleright Obtain the decomposed matrices of $W$ from $\text{Set}_{SVD}$
15:              $U^{\prime}=WX^{\prime}X^{\prime T}(S^{-1})^{T}V\Sigma_{1}(\Sigma_{1}V^{T}S^{-1}X^{\prime}X^{\prime T}(S^{-1})^{T}V\Sigma_{1})^{-1}$triangleright Closed-form
16:              $W_{u}\leftarrow U^{\prime}(\Sigma_{1})^{1/2},W_{v}\leftarrow S^{-1}V(\Sigma_{1})^{1/2}$triangleright Obtain two low-rank matrices
17:              $L(W)\leftarrow L(W_{u},W_{v})$triangleright Replace $W$ with $W_{u}$ and $W_{v}$ in $L$
18:         end for
19:         $X^{\prime}\leftarrow L(X^{\prime})$triangleright Use the compressed layer to calculate the new input activation $X^{\prime}$ for the next layer
20:     end for
21:     return $M^{\prime}$
22:end procedure
Table 9: Perplexity of LLaMA-7B compressed by SVD-LLM and ASVD (w/ and w/o LoRA) on WikiText-2.
Method 20% 30% 40% 50% 60%
ASVD 11.14 51 1407 15358 57057
ASVD + LoRA 7.37 10.16 14.86 21.83 44.81
SVD-LLM 7.94 9.56 13.11 23.97 53.74
SVD-LLM + LoRA 8.28 9.14 10.65 13.26 17.93
A.3 Performance with LoRA Fine-Tuning.
LoRA [13] is a common fine-tuning technique for LLM. It has been applied with pruning-based LLM compression methods such as LLM-Pruner [16] to mitigate accuracy drop after pruning. LoRA can also be combined with SVD-based LLM compression methods by modifying the forward pass of a linear layer as:
$Y$ $=W_{u}^{\prime}W_{v}^{\prime}X$ (8)
To examine the performance of SVD-LLM in combination with LoRA, we follow the same configuration used in LLM-Pruner [16] to fine-tune LLaMA-7B compressed by SVD-LLM and ASVD under the compression ratios from 20% to 80% with LoRA. The results are shown in Section A.2. We have three observations. (1) Comparing SVD-LLM with SVD-LLM $+$ LoRA, under the compression ratio between 20% and 50%, the accuracy enhancement brought by LoRA is limited; as the compression ratio increases, LoRA plays a more significant role in improving the performance of SVD-LLM. (2) Compared to ASVD $+$ LoRA, SVD-LLM $+$ LoRA consistently achieves better accuracy across all the compression ratios. In particular, under 70% compression ratio, the perplexity of SVD-LLM $+$ LoRA is 75% lower than that of ASVD $+$ LoRA. (3) Finally, even without LoRA, SVD-LLM is able to achieve perplexity comparable to ASVD $+$ LoRA, especially under the compression ratios between 20% and 60%.
A.4 Generated Content from the Compressed Model
We also compare some examples of sentences generated by LLaMA-7B compressed with SVD-LLM and ASVD in Table 10. As shown, the sentences generated by the model compressed by SVD-LLM exhibit better fluency, relevance, and informativeness compared to that compressed by ASVD. More importantly, when the compression ratio is increased to 30%, the previous state-of-the-art method ASVD completely loses its generation ability. In contrast, even when the compression ratio is up to 40%, SVD-LLM is still capable of generating complete sentences.
Table 10: An example of contents generated by the compressed LLaMA-7B by SVD-LLM and and under different compression ratios. The input is marked in bold and the normal texts are the generated sentences.
Ratio SVD-LLM ASVD
Original Building a website can be done in 10 simple steps: Step 1: Register an domain name; Step 2: Get a hosting package; Step 3: Develop a website; Step 4: Design the website; Step 5: Make the website live; Step 6: Perform on-page SEO; Step 7: Promote the website; Step 8: Evaluate the results; Step 9: Improve the website; Step 10: Build a website repeat …
20% Building a website can be done in 10 simple steps: 1) Identifying an appropriate theme. 2 ) Determining which types of Web site would be of interest and which could be implemented with minimal overhead and resources. 3 ) Identify the audience 4 ) Choose Web site software. 5 ) Select Web site developers. 6 ) Design Web site 7 ) Identify Web site technology 8 ) Design Web site layout 9 ) Define Web site aesthetics and 10 ) Define Web site access controls … Building a website can be done in 10 simple steps: 1, Build up your website. You will need to select a name for your website. As a beginner, it is a good idea to select your company name or product name as your website name. You can use your real company or product name or you can create an unusual business name. This will make the name of your business name consistent throughout your website. 2, Decide on your business name. You will be choosing the official name of your business for all activities that represent your business. This includes your business name, products, your website, and any …
30% Building a website can be done in 10 simple steps: 1. Preliminaries and planning 2. Designing 3. Prototyping 4. Production 5. Operation… Building a website can be done in 10 simple steps: To: With A Web You Are Into 2 Zim 40 Wh I M T If Y As An Onl You Orwice There W Os You Hcimp …
40% Building a website can be done in 10 simple steps: 2 days of website plan. If you want to set a web site. This is the 1st step. In the 2 days . It wants the company to develop a website to help the investors and investors. He will have another investor. He wants to help others. The company is working to grow, and create a name for some investors and investors. 100 million US dollars It is used to build a hotel and a hotel. In order to obtain their cooperation, investors, investors, investors in . Building a website can be done in 10 simple steps: Front 1222ED825333 $3333$etennenn eenneenneSecondinjustanyanojunatajin …
A.5 More Experiments on compressing different LLMs
Table 11: Zero-shot performance comparison of OPT-6.7B, LLaMA 2-7B, Vicuna-7B, Mistral-7B compressed by SVD-LLM and baselines under 20% compression ratio on seven common sense reasoning datasets (measured by both individual and average accuracy ($\uparrow$)). The best performance is marked in bold. The relative performance gain compared to the best-performing baseline is marked in green color inside bracket.
[…]
OPT-6.7B SVD 0.14 0.27 0.51 0.25 0.23 0.54 0.21 0.31
FWSVD 0.15 0.26 0.49 0.26 0.20 0.52 0.21 0.30
ASVD 0.14 0.40 0.51 0.30 0.20 0.59 0.22 0.34
SVD-LLM 0.24 0.60 0.60 0.45 0.28 0.73 0.24 0.45 ($\uparrow$32%)
LLaMA 2-7B SVD 0.15 0.27 0.49 0.26 0.23 0.52 0.20 0.30
FWSVD 0.12 0.25 0.49 0.25 0.22 0.52 0.21 0.30
ASVD 0.25 0.31 0.60 0.41 0.32 0.72 0.23 0.41
SVD-LLM 0.26 0.50 0.60 0.41 0.26 0.66 0.23 0.41 ($\uparrow$0%)
Mistral-7B SVD 0.14 0.25 0.52 0.26 0.23 0.54 0.20 0.30
FWSVD 0.15 0.28 0.52 0.26 0.21 0.53 0.21 0.31
ASVD 0.21 0.51 0.58 0.42 0.25 0.50 0.26 0.39
SVD-LLM 0.17 0.55 0.58 0.36 0.25 0.67 0.21 0.40 ($\uparrow$3%)
Vicuna-7B SVD 0.15 0.26 0.50 0.26 0.23 0.52 0.20 0.30
FWSVD 0.14 0.27 0.49 0.26 0.22 0.53 0.20 0.30
ASVD 0.21 0.53 0.55 0.39 0.30 0.41 0.23 0.37
SVD-LLM 0.23 0.51 0.58 0.40 0.28 0.67 0.22 0.41 ($\uparrow$11%)
We also evaluate the performance of different LLMs, including OPT-6.7B, LLaMA 2-7B, Mistral-7B, and Vicuna-7B under 20% compression ratio on seven common sense reasoning datasets. The results are shown in Section A.5. SVD-LLM performs better than the best-performing baseline in most of the datasets across different LLMs and even achieves 32% higher average accuracy on OPT-6.7B.